Which quadratic equation has solutions of 5 and 7?
Answer by ewatrrr(24785)

 
Hi
Which quadratic equation has solutions of 5 and 7?
The factor theorem states that a polynomial f(x) has a factor (x − k)
if and only if f(k) = 0
(x-5)(x-7) = x^2 - 12x + 35